OpenStudy (anonymous):

Find the number of paid hours worked from 8:01 a.m. to 4:00 p.m., assuming pay is for full half-hours.

OpenStudy (anonymous):

I know work this in my head but they want me to show this on paper.

OpenStudy (anonymous):

\[\text{Mod}[60(16-8)-1,30]*\frac{1}{2}=\frac{29}{2}=14.5 \text{ hours} \]

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

You are welcome.